Announcement

Collapse
No announcement yet.

Help on Step 6 Please

Collapse
X
 
  • Filter
  • Time
  • Show
Clear All
new posts

  • Help on Step 6 Please

    Hello,
    I've been working on my Minipulse kit, taking my time and going step by step. I've been using a friend's oscilloscope to test along the way. Everything has been looking good until I got to step 6. When testing I am getting inconsistent results. TP5 and TP6 seem OK, but TP7 is strange. If I hook the scope up, to TP5 on one channel and TP7 on the other, when I first power it up, it looks OK, but after a few seconds the pulse width of the main sample delay seems to become unstable and then everything on TB7 goes dead. If I wait a while and come back, I can recreate the symptoms. It's almost like it is able to reset itself after a while and then it craps out again. I went back through all the steps performing each test and everything else looks fine. I attached a link to some pics of my board as it is and one scope shot of when TP5 (channel 2) and TP7 (channel 1) both appear to be working.The only thing I've noticed that is different is that the parts list calls for a pair of MC14538BCPs, but my kit has HCF4538BEs. I looked at the spec sheets of both, and from what I can tell, they are very similar. I went ahead and ordered a pair of the MC14538BCPs to try. I also tried swapping the two HCF4538BEs, with each other, no change. I'm losing my mind.What else can I test? Have I done something obviously wrong?

    https://goo.gl/photos/BGF8LN8JwnkGMope8

    Here's also a short video of TP7 going dead.
    https://goo.gl/photos/yHDY1rAzVGe49qRd9

    I appreciate any help.
    Thanks,
    Chris

  • #2
    In your video the scope appears to be triggered from TP7. When the signal on TP7 disappears, the scope loses sync and the main sample pulse width (TP5) starts hopping around the display. So, either the pulse at TP7 is dropping out, or it's a problem with the scope channel connected to TP7.

    Firstly I would swap the scope channels over, and then see if the fault on TP7 remains, or whether it shifts to TP5.

    If not, then there's a problem within the pulse generator section. Check what's going on with TP8 (EFE sample pulse). Is this ok, or does it also disappear after the same period of time?

    Comment


    • #3
      Thanks for the reply.
      You're correct about the trigger. Later in the video I switched the trigger to the other channel to make it stop hopping. I didn't check TP8 at that time, but the night before, it seemed to be working. I have to admit though, I first came across this issue at my friend's house when I was using his scope. I came home and reflowed all of the solder joints, made sure all the flux was cleaned off. I borrowed his scope and tested again when I made the video. Later I realized that I had forgotten to short the pins on TP3! I'll retest tonight and make sure to have the pins shorted. I'll check TP7 and TP8 and post a new video.

      Comment


      • #4
        Originally posted by Calla969 View Post
        I also tried swapping the two HCF4538BEs, with each other, no change. I'm losing my mind.What else can I test? Have I done something obviously wrong?
        if you swapped the ICs (U6,U7) with each other, and there is NO problem at TP5, then the ICs are OK.
        The problem is in the connections of IC U7, check it opticaly that all pins are ok to the right place, and check the R11,C11,D7,D4.

        Solder again the U7, i see on the picture, that pin4 has less solder ,or not soldered, ...and press the ic correct to the socket.

        Comment


        • #5
          I'll check it again and add some solder. Thanks.

          Comment


          • #6
            Alright, I added some solder to U7, I really think it was fine, just looked light in the picture. I remembered to short pins 1 and 2 on PL3. When I hooked up TP5 and TP7, I'm getting similar results. I visually inspected everything and checked the values of the resistors. I tried to verify the value of the capacitor, but on my cheapo tester, caps measure as diodes if I measure them on the board. I measured each component before installing though. C11 calls for a 2n2, my kit came with a 222J100 which is what I installed at C11. TP5 and TP6 seem OK. TP7 and TP8 seem all over the place. I linked some scope shots.

            The one with the nice clean square waves is TP5 on ch 1 and TP6 on ch 2. The messy looking one is TP7 on ch 1 and TP8 on CH 2.

            https://goo.gl/photos/uFRmDcAiJfLcHLWc6

            Comment


            • #7
              I am also using HCF4538BEs, so that's not the problem.

              As far as I can see from the PCB image, you have the correct component values fitted. The two diodes can be checked easily using the diode checker on a multimeter, even while they are in circuit.

              From the image of TP5/TP6, you seem to have the scope set to 20V/div AC. Firstly, change the scope input coupling to DC, and check that your scope probes are not set to x10. They should be on x1.
              If we assume the probes are [incorrectly] set to x10, then the correct amplitude is 5V. BUT - the pulse should be going from 0V to -5V, not +5V.
              Where is your ground probe connected? Remember that the MPP uses the battery positive as 0V, and the ground probe should be on TP10.

              From the image of TP7/TP8, the amplitude is only +1.4V, which should (of course) be negative. Also (unless you switched the probes onto x1) the actual voltage is 140mV.

              Lastly, I assume your scope's timebase setting is on 50us/dev. In which case the pulse widths seem correct.

              Comment


              • #8
                Thanks for the help! I'll check this all out tonight. I am pretty clueless when it comes to using a scope. This is the first project I've done where I needed one. I'm sure I haven't been using the correct settings. Maybe everything is actually fine.

                Comment


                • #9
                  Last night I connected the scope, correctly this time, and checked pins 4 - 8. I think all look OK except for 7 and 8....still. At first, they looked alright, but that only lasted a few seconds and then they both went completely dead. I swapped out the IC with a spare and that didn't change anything. I connected the scope with my 12v power supply turned off. When I turn the power on, I see some activity on 7 and 8 for just a second, then it goes back to what it looked like with no power. Dead. What could be causing this? The resistors and diodes seem fine. Should I replace all of the caps that I installed in step 6? What about the 100k trimmer? It appears to be functioning correctly.

                  Comment


                  • #10
                    Originally posted by Calla969 View Post
                    Last night I connected the scope, correctly this time, and checked pins 4 - 8. I think all look OK except for 7 and 8....still. At first, they looked alright, but that only lasted a few seconds and then they both went completely dead. I swapped out the IC with a spare and that didn't change anything. I connected the scope with my 12v power supply turned off. When I turn the power on, I see some activity on 7 and 8 for just a second, then it goes back to what it looked like with no power. Dead. What could be causing this? The resistors and diodes seem fine. Should I replace all of the caps that I installed in step 6? What about the 100k trimmer? It appears to be functioning correctly.
                    Can you capture and post the scope images for TP5 to TP8, as before (but this time with the correct scope settings)?
                    Please don't randomly replace components. That's not the correct way to debug the circuit.

                    Comment


                    • #11
                      For both, CH1 is red, CH2 is yellow. Probes are set to 1x, grounded on TP10, DC coupling.

                      These 2 pictures are with CH1 on TP6 and CH2 on TP5. Both channels on 20v/div and 20us/div. Trigger is on CH1 at -5.6v. Both pics are the same, I just separated them on one image so they can be seen easier.
                      https://goo.gl/photos/ba9dDmCt6zSeaAP96

                      Here's a video with CH1 on TP8 and CH2 on TP7. It starts out on 5v/div, but I switch it to 10v/div while it's going crazy trying to be able to see more of what's going on. I have to act fast because it goes completely flat pretty quickly and I only seem to even get this reaction if I let it sit for 20 minutes or so. Trigger is on CH1.
                      https://goo.gl/photos/62Bp6EQp55djjsuB7

                      Comment


                      • #12
                        Originally posted by Calla969 View Post
                        For both, CH1 is red, CH2 is yellow. Probes are set to 1x, grounded on TP10, DC coupling.

                        These 2 pictures are with CH1 on TP6 and CH2 on TP5. Both channels on 20v/div and 20us/div. Trigger is on CH1 at -5.6v. Both pics are the same, I just separated them on one image so they can be seen easier.
                        https://goo.gl/photos/ba9dDmCt6zSeaAP96
                        Assuming you have correctly switched the probes to x1, then there must be something wrong with the scope settings ... since there is no way channels 1 and 2 should be on 20V/div. If this were true, the amplitude of the signals shown in the photo would be -50V. Presumably there is a x10 setting somewhere on the scope itself. If that is the case, then the signal amplitude is really -5V, which is correct. Please recheck your scope settings.

                        Originally posted by Calla969 View Post
                        Here's a video with CH1 on TP8 and CH2 on TP7. It starts out on 5v/div, but I switch it to 10v/div while it's going crazy trying to be able to see more of what's going on. I have to act fast because it goes completely flat pretty quickly and I only seem to even get this reaction if I let it sit for 20 minutes or so. Trigger is on CH1.
                        https://goo.gl/photos/62Bp6EQp55djjsuB7
                        Unless you've changed something between the first test and this one, your 5V/div setting is most likely to really be 500mV/div. Unless the signal amplitude is close to -5V, it's probably not there at all, and what you're seeing is some low level noise.

                        OK - let's see if we can progress this further:
                        Remove U7 (HCF4538BE) from its socket, and carefully bend pins 4 and 10 out to the side. Then put the IC back into the socket with these 2 pins hanging over the side of the socket. What I'm looking for here is whether something is loading the signals on TP7 and TP8. Repeat the previous tests and see whether there is now a valid signal on U7 pins 4 and 10 (Obviously there will not be a signal on TP7 or TP8 at this point, since these have been disconnected.) If so, then I suspect there's a problem with Q3 and Q4.

                        Comment


                        • #13
                          Alright, last night I came home from work with a debilitating headache, so I didn't spend very much time with this, but I did find the setting in the scope and it was set to 10x. Probes were 1x, scope was 10x. I set the scope to 1x and retested TP7 and TP8. It looks much better now. I played around with it for a few minutes and it seemed stable. I attached an image below. I think you can see the settings in the image, 5v/div, 20us/div. I didn't note which pin was on which channel, I could barely even see what I was doing due to the headache. I didn't bend the pins of U7 since I was getting stable results.
                          https://goo.gl/photos/qhi9pH1Xqj67G2gk8

                          Here's TP5 and TP6.
                          https://goo.gl/photos/KURj3Ydyz1r7f7io7

                          If this looks decent, I'll move ahead and test TP4 from step 7. Hopefully that will be OK and I can get this thing finished. I really appreciate the help.

                          Comment


                          • #14
                            Originally posted by Calla969 View Post
                            Alright, last night I came home from work with a debilitating headache, so I didn't spend very much time with this, but I did find the setting in the scope and it was set to 10x. Probes were 1x, scope was 10x. I set the scope to 1x and retested TP7 and TP8. It looks much better now. I played around with it for a few minutes and it seemed stable. I attached an image below. I think you can see the settings in the image, 5v/div, 20us/div. I didn't note which pin was on which channel, I could barely even see what I was doing due to the headache. I didn't bend the pins of U7 since I was getting stable results.
                            https://goo.gl/photos/qhi9pH1Xqj67G2gk8

                            Here's TP5 and TP6.
                            https://goo.gl/photos/KURj3Ydyz1r7f7io7

                            If this looks decent, I'll move ahead and test TP4 from step 7. Hopefully that will be OK and I can get this thing finished. I really appreciate the help.
                            If you look at the shape of the signals on TP5 and TP6, you will notice that they swing from 0V to -5V, and are very clean. Both TP7 and TP8 should also have the same amplitude and be just as clean, but your scope image does not show this to be the case. So there's still something wrong somewhere. However, the pulses widths look correct at ~58us.

                            Ignore what I said earlier about Q3 and Q4, as you don't appear to have fitted these just yet. If we assume that nothing is loading the outputs of U7 (since at the moment there are no other components connected there) then the problem must be located at the input of U7. In that case, can you post a scope image showing the signals on U7 pins 2 and 14?

                            It would also be interesting to see the signals at U7 pins 7 and 9.

                            Comment


                            • #15
                              Actually, Q3 and Q4 are installed. At one point, I thought I was OK on TP7 and TP8 and I moved on to step 7, just haven't tested it yet with the correct scope settings, but when I tested it with a volt meter, TP4 was reading around 2.5v instead of 500mv. I did adjust R6 to what I believe is the correct 22us delay. I'll post relevent scope images tonight. Again, thanks for all the help. I'm learning a lot with this project, that was the goal.

                              Comment

                              Working...
                              X